# Condensate Size Control by Charge Asymmetry

### v1

@@ Abstract
To test whether our predictions hold more generally, and in particular in 3D systems, we return to our MD simulations (details described in the Appendix). Fig. 4a shows that larger charge asymmetry $\Delta z$ indeed leads to smaller droplets. We also observe system spanning droplets close to charge neutrality ($\Delta z=0$; yellow symbols), which are consistent with macrophase separation (region H+H in Fig. 3). The detailed size distributions shown in Fig. S2 in the Appendixindicate this dependence unambiguously, although finite size effects are clearly visible for small charge asymmetries. Moreover, Fig. 4b shows that varying the interaction strength $\epsilon_{\mathrm{LJ}}$ between the polyions affects the droplet size only weakly, consistent with the weak influence of $\chi$ shown in Fig. 3d. In both cases, at weak charge asymmetry droplet size slightly decreases for larger attraction, conversely to the behavior at strong charge asymmetry. Taken together, these data confirm the predictions from the field theory, suggesting that droplet size control by net charges is a general phenomenon.
@@
The described droplet size control essentially emerges from a trade-off between short-ranged attraction driving phase separation and long-ranged electrostatic repulsion if droplets accumulate net charges. At weak charge asymmetry, the system exhibits macrophase separation since the attraction dominates. Conversely, for strong charge asymmetry, the system will stay homogeneous to avoid accumulating net charges. Between these two extremes, the patterned phases emerge as a compromise between phase separation and accumulating net charge. This behavior is reminiscent of phase separation in elastic media, where elastic effects mediate long-ranged repulsion, resulting in patterned phases and a continuous transition to the homogeneous state [25]. This similarity suggests that droplet size can be generally controlled when the short-ranged attraction leading to phase separation is opposed by some long-ranged repulsion, reminiscent of classical pattern formation [27]. Another example for this mechanism are chemically active droplets, where long-ranged repulsion emerges from a reaction-diffusion system [28, 29, 24, 30], which can in fact be interpreted using an electrostatic analogy [31, 32, 33].
@@
Here, we focused on the most basic mechanism of droplet size control by net charges. For instance, we assumed that polyions have a fixed charge, whereas realistic molecules will exhibit charge regulation, so their behavior will depend on pH. Molecules involved in complex coacervation, and particularly in biomolecular condensates, are also much more complex: their charge distribution will be inhomogeneous, the unspecific interactions will be heterogeneous and might depend on charge, and the polymer structure cannot be ignored, particularly for nucleic acids and intrinsically disordered protein regions [34]. It will be interesting to see how such effects influence the described size control and how these charge asymmetries affect condensates inside cells. It is already tempting to speculate that charged droplets were implicated in the origin of life [35, 36], so that their size is maintained and the associated Rayleigh instability [18, 17] could explain division.
Acknowledgements
@@
We thank Jan Kirschbaum for stimulating discussions, and we thank Nynke Hettema, Noah Ziethen and Athanassios Panagiotopoulos for helpful comments on our manuscript. We gratefully acknowledge funding from the Max Planck Society and the European Union (ERC, EmulSim, 101044662), the Chan Zuckerberg Initiative DAF (an advised fund of Silicon Valley Community Foundation; grant 2023-332391 ), and departmental start-up funds via the Department of Chemical and Biological Engineering and the Omenn–Darling Bioengineering Institute at Princeton University. This research was partially supported by the National Science Foundation (NSF) through the Princeton University (PCCM) Materials Research Science and Engineering Center DMR-2011750.
@@ A.1 Simulation procedure
Molecular dynamics simulations are run in the $NVT$ ensemble with neutral mixtures of polyions and counterions in implicit solvent (Fig. 1) using Lennard-Jones units. All particles are modeled with a size $\sigma$ and a mass $m$. The total energy of the simulation system is given by:
$U=U_{\mathrm{el}}+U_{\mathrm{CA}}+U_{\mathrm{EV}},$ (S1)
where $U_{\mathrm{el}}$ describes the long-range electrostatic interactions between charged monomers, $U_{\mathrm{CA}}$ describes the same chain bonded interactions and short-range heterotypic attraction between polycations $C$ and polyanions $A$, and $U_{\mathrm{EV}}$ describes the ion-ion, ion-polymer, and homotypic polymer-polymer excluded-volume interactions. More specifically, $U_{\mathrm{el}}$ is modeled using Coulomb’s potential:
$U_{\mathrm{el}}=\sum_{\alpha,\beta}\sum_{i,j}\frac{\alpha_{i}\beta_{j}}{\varepsilon r_{ij}}.$ (S2)
In this form, $\alpha,\beta\in\{e^{+},e^{-},z_{\mathrm{C}},z_{\mathrm{A}}\}$ describes the type of the charged monomers $i$ and $j$, $r_{ij}$ is the distance between the charged monomers, and the dielectric constant is $\varepsilon$. Next, $U_{\mathrm{CA}}$ models the combination of the bonded potential $V_{\mathrm{bond}}$ and the Lennard-Jones potential $V_{\mathrm{LJ}}$ with cut-off and shifting:
$U_{\mathrm{CA}}=\sum_{i=1}^{N}\sum_{j=1}^{n-1}V_{\mathrm{bond}}(r_{j+1,j})+\sum_{i,j}\left(V_{\mathrm{LJ}}(r_{i,j})-V_{\mathrm{LJ}}(r_{\mathrm{c}})\right).$ (S3)
Here, $N$ is total number of polymers in the system, $n$ is the polymer length. The bonded potential is modeled using a harmonic potential $V_{\mathrm{bond}}(r)=K(r-r_{b})^{2}$. As for the Lennard-Jones potential we use:
@@
Finally, for the excluded volume for all the particles (besides interactions between $C$ and $A$), we use the Weeks-Chandler-Andersen (WCA) potential
@@
In all simulations, we use $N=512+512=1024$ polyions and each polyion is composed of $n=10$ bonded monomers. We use a bond coefficient of $100.0\epsilon_{\mathrm{LJ}}$ and an equilibrium distance $r_{\mathrm{b}}=1.0\sigma$. For the Lennard-Jones interactions, we use a cut-off of $r_{\mathrm{c}}=2.5\sigma$ and attraction strength of $\epsilon_{\mathrm{LJ}}=1.0$ which is also the energy scale for other interactions. The charge of polyions are varied by simulation. The charge of counterions are set to have a magnitude of $1.0$. Electrostatic interactions are calculated using a PPPM grid for interactions beyond $5.0\sigma$. Electrostatic interactions within $5.0\sigma$ are calculated directly. The dielectric constant is set to $\varepsilon=80.0$. The volume fractions of each species in the simulation box are set to be $\phi_{\mathrm{C}}=\phi_{\mathrm{A}}=0.1$ and $\phi_{e^{-}}=z_{\mathrm{C}}\phi_{\mathrm{C}}$ and $\phi_{e^{+}}=-z_{\mathrm{A}}\phi_{\mathrm{A}}$.
Each simulation trajectory is first annealed at a temperature of $k_{\mathrm{B}}T/\epsilon_{\mathrm{LJ}}=5.0$ for $500\tau$ using a Langevin thermostat with a damping of $100.0\tau$. Thereafter, the temperature of each simulation is set to $k_{\mathrm{B}}T/\epsilon_{\mathrm{LJ}}=1.0$ also using a Langevin thermostat with a damping of $100.0\tau$. Simulation trajectories are then run for $110$ million $\tau$. The last $10$ million $\tau$ is collected for data analysis, sampling a frame every $10,000\tau$. We use a timestep of $0.05\tau$ in the simulations.
Further, in the analysis of MD simulations, droplets are defined using a cutoff of $2.0\sigma$ between polyions. The trajectory analysis software Ovito [37] is used to generate cluster data from simulation frames. An in-house code is used to generate radial density profiles from an Ovito [37] cluster analysis pipeline.
@@ A.2 Additional result figures
Additional figures are presented in this section corresponding to results from MD simulations. Figure S1 presents radial charge density profiles from MD simulations at ($z_{\mathrm{C}}=0.55$, $z_{\mathrm{A}}=-0.55$), ($z_{\mathrm{C}}=0.64$, $z_{\mathrm{A}}=-0.3$), and ($z_{\mathrm{C}}=1.05$, $z_{\mathrm{A}}=-0.34$). See Fig. 1 for a visualization of these droplets. Figure S2 shows normalized droplet size distributions corresponding to the four data points in Fig. 4a. Figure S3 confirms that the relationship between charge asymmetry and droplet size holds across three different values of $\epsilon_{\mathrm{LJ}}$. Finally, Figure S4 quantifies the standard deviations in normalized droplet sizes for the simulations presented in Fig. 4a.
@@
Figure S1: Radial charge density profiles indicate charge asymmetry controls droplet size. Radial charge density profiles (RDPs) for (a) $z_{\mathrm{C}}=0.55$, $z_{\mathrm{A}}=-0.55$ (b) $z_{\mathrm{C}}=0.64$, $z_{\mathrm{A}}=-0.3$ and (c) $z_{\mathrm{C}}=1.05$, $z_{\mathrm{A}}=-0.34$. Shaded regions indicate plus and minus the standard deviation from three equal blocks of the 1,000 frames used in generating the RDPs. The profiles indicate that increasing charge asymmetry decreases droplet size. See Fig. 1 for visualization of these droplets.
@@
Figure S2: Larger charge asymmetries create broader distributions of droplet size. Four simulation systems, (I) $z_{\mathrm{C}}=0.89$, $z_{\mathrm{A}}=-0.54$ (II) $z_{\mathrm{C}}=0.99$, $z_{\mathrm{A}}=-0.5$ (III) $z_{\mathrm{C}}=1.09$, $z_{\mathrm{A}}=-0.46$ (IV) $z_{\mathrm{C}}=1.19$, $z_{\mathrm{A}}=-0.42$, corresponding to Fig. 4a are plotted in histograms of normalized droplet sizes. At larger charge asymmetries (III and IV), the distribution of cluster sizes is larger and the average droplet size is smaller. Finite-sized effects prevail at smaller charge asymmetries (I and II) indicated by the multi-modal distributions (i.e., size discrepancy among sustained droplets).
@@
Figure S3: Charge asymmetry’s effect on droplet size is preserved at higher polyion attraction strengths. The normalized droplet size for four charge pairings (($z_{\mathrm{C}}=0.89$, $z_{\mathrm{A}}=-0.54$), ($z_{\mathrm{C}}=0.99$, $z_{\mathrm{A}}=-0.5$), ($z_{\mathrm{C}}=1.09$, $z_{\mathrm{A}}=-0.46$), ($z_{\mathrm{C}}=1.19$, $z_{\mathrm{A}}=-0.42$) corresponding to I-IV in Fig. 4a) at three different attraction strengths between oppositely charged polyions ($\epsilon_{\mathrm{LJ}}=1.0,1.25,1.5$). As charge asymmetry is increased the normalized droplet size in the simulations decreases regardless of the value of $\epsilon_{\mathrm{LJ}}$. Error bars represent standard deviation in the droplet size data at every condition.

### v2

@@ Abstract
To test whether our predictions hold more generally, and in particular in 3D systems, we return to our MD simulations (details described in the Supporting Text [20]). Fig. 4a shows that larger charge asymmetry $\Delta z$ indeed leads to smaller droplets. We also observe system spanning droplets close to charge neutrality ($\Delta z=0$; yellow symbols), which are consistent with macrophase separation (region H+H in Fig. 3). The detailed size distributions shown in Supporting Figure S2 [20] indicate this dependence unambiguously, although finite size effects are clearly visible for small charge asymmetries. Moreover, Fig. 4b shows that varying the interaction strength $\epsilon_{\mathrm{LJ}}$ between the polymers affects the droplet size only weakly, consistent with the weak influence of $\chi$ shown in Fig. 3d. In both cases, at weak charge asymmetry droplet size slightly decreases for larger attraction, conversely to the behavior at strong charge asymmetry. Overall, these data confirm the predictions from the field theory, suggesting that droplet size control by net charges is a general phenomenon.
One surprising feature of our model is the fact that salt, i.e., excess ions that are not necessary to neutralize the polymers, cannot screen the electrostatic interactions. The concentration profiles obtained from MD simulations (Fig. S1 [20]) and the field theory (Fig. 2) suggest that the interactions between polymers prevent ions from moving into the droplet to neutralize it, thus building up a net charge. Consequently, the system cannot be described as a dilute electrolyte with mobile ions, e.g., using Debye–Hückel theory [9]. In particular, the intuitive argument that size-control would be limited by the Debye screening length is thus misleading. Instead, ions need to be treated explicitly, which is likely a general feature of condensates involving complex, charged molecules [28, 29] and similar systems [30, 31, 32].
@@
The described droplet size control essentially emerges from a trade-off between short-ranged attraction driving phase separation and long-ranged electrostatic repulsion if droplets accumulate net charges. At weak charge asymmetry, the system exhibits macrophase separation since the attraction dominates. Conversely, for strong charge asymmetry, the system will stay homogeneous to avoid accumulating net charges. Between these two extremes, the patterned phases emerge as a compromise between phase separation and accumulating net charge. Such patterned states also emerge in simpler phase separating system with Coulomb-like interactions [33, 34, 35, 36, 37]. In particular, when elastic effects mediate long-ranged repulsion, the patterned phase also exhibits a continuous transition to the homogeneous state [25]. This similarity suggests that droplet size can be generally controlled when the short-ranged attraction leading to phase separation is opposed by some long-ranged repulsion [38], reminiscent of classical pattern formation [39]. Another example for this mechanism are chemically active droplets, where long-ranged repulsion emerges from a reaction–diffusion system [40, 41, 24, 42], which can in fact be interpreted using an electrostatic analogy [43, 44, 45].
@@
Here, we focused on the most basic mechanism of droplet size control by net charges. For instance, we assumed that polymers have a fixed charge, whereas realistic molecules will exhibit charge regulation, so their behavior will depend on pH. Molecules involved in typical biomolecular condensates are also much more complex: their charge distribution will be inhomogeneous, the nonspecific interactions will be heterogeneous and might depend on charge, and the polymer structure cannot be ignored, particularly for nucleic acids and intrinsically disordered protein regions [46]. Such effects inform short-ranged interactions within biomolecular condensates [1, 3], but the impact on long-ranged electrostatic is less understood. Our work shows that such long-ranged interactions can have profound consequences that cannot be explained by the simple Debye–Hückel theory. Therefore, it will be interesting to see how the described size control and ion distribution behaves in more complex systems as well as how these charge asymmetries affect condensates inside cells. It is already tempting to speculate that charged droplets were implicated in the origin of life [47, 48], so that their size is maintained and the associated Rayleigh instability [16, 15] could even explain division.
Acknowledgements.
@@
We thank Jan Kirschbaum for stimulating discussions, and we thank Nynke Hettema, Noah Ziethen and Athanassios Panagiotopoulos for helpful comments on our manuscript. We gratefully acknowledge funding from the Max Planck Society and the European Union (ERC, EmulSim, 101044662), the Chan Zuckerberg Initiative DAF (an advised fund of Silicon Valley Community Foundation; grant 2023-332391 ), the National Institute of General Medical Sciences of the National Institutes of Health under Award Number R35GM155259, and departmental start-up funds via the Department of Chemical and Biological Engineering and the Omenn–Darling Bioengineering Institute at Princeton University. This research was partially supported by the National Science Foundation (NSF) through the Princeton University (PCCM) Materials Research Science and Engineering Center DMR-2011750. The simulations presented in this article were performed on computational resources managed and supported by Princeton Research Computing, a consortium of groups including the Princeton Institute for Computational Science and Engineering (PICSciE) and the Office of Information Technology’s High Performance Computing Center.
@@ Appendix A Details of the molecular dynamics (MD) simulations
Here we present additional details on the MD simulations performed in this investigation. In particular, we provide additional details on the simulation procedure in subsection A.1. Additional figures pertaining to the data presented in the main text are presented in subsection A.2. In subsection A.3 we show data that demonstrates the simulations discussed in the main text are sampling equilibrium statistics. Further, we explore the impact of ion size on the mechanism of droplet size control by investigating a subset of simulations with reduced ion sizes in subsection A.4. Finally, we validate that the damping constant of the Langevin thermostat does not alter simulation results if simulated in the overdamped limit (as opposed to the larger damping constant used in the simulations for the main text) in subsection A.5.
@@ A.1 Simulation procedure
All simulations are performed in the MD software LAMMPS [49]. MD simulations are run in the $NVT$ ensemble with neutral mixtures of charged polymers and ions in implicit solvent (Main Figure 1a). The equations of motion are integrated using the velocity-Verlet algorithm and the temperature is controlled using a Langevin thermostat (discussed in more detail below). All particles (i.e., monomers and ions) are modeled with a unit size $\sigma$ and a unit mass $m$. Note that all simulation quantities reported in this SI and in the main text are in reduced Lennard–Jones units.
The total energy of the simulation system is given by:
$U=U_{\mathrm{el}}+U_{\mathrm{P}}+U_{\mathrm{EV}},$ (S1)
where $U_{\mathrm{el}}$ describes the long-range electrostatic interactions between charged species (i.e., charged monomers on the polymers as well as ions), $U_{\mathrm{P}}$ describes the same chain bonded interactions and short-range heterotypic attraction between positively charged monomers $P^{+}$ and negatively charged monomers $P^{-}$, and $U_{\mathrm{EV}}$ describes the ion–ion, ion–polymer, and homotypic polymer–polymer excluded-volume interactions. More specifically, $U_{\mathrm{el}}$ is modeled using Coulomb’s potential:
$U_{\mathrm{el}}=\sum_{i,j}\frac{z_{i}z_{j}}{\varepsilon r_{ij}}.$ (S2)
In this form, $z_{i}$ and $z_{j}$ describe the charge on species $i$ and $j$, $r_{ij}$ is the distance between the charged species, and the dielectric constant is $\varepsilon$. Next, $U_{\mathrm{P}}$ models the combination of the bonded potential $V_{\mathrm{bond}}$ and the Lennard-Jones potential $V_{\mathrm{LJ}}$ with cut-off and shifting:
$U_{\mathrm{P}}=\sum_{i=1}^{N}\sum_{j=1}^{n-1}V_{\mathrm{bond}}(r_{j+1,j})+\sum_{i,j}\left(V_{\mathrm{LJ}}(r_{i,j})-V_{\mathrm{LJ}}(r_{\mathrm{c}})\right).$ (S3)
Here, $N$ is total number of polymers in the system, $n$ is the polymer length. The bonded potential is modeled using a harmonic potential $V_{\mathrm{bond}}(r)=K(r-r_{\text{b}})^{2}$. As for the Lennard-Jones potential we use:
@@
Finally, for the excluded volume for all the particles (besides interactions between $P^{+}$ and $P^{-}$), we use the Weeks–Chandler–Andersen (WCA) potential
@@
In all simulations, we use a box size of $60\times 60\times 60\>\sigma^{3}$ with our species volume fractions of $\phi_{P^{+}}=\phi_{P^{-}}=0.1$ and $\phi_{e^{-}}=z_{P^{+}}\phi_{P^{+}}$ and $\phi_{e^{+}}=-z_{P^{-}}\phi_{P^{-}}$. This corresponds to $N=512+512=1024$ polymers where each polymer is composed of $n=10$ bonded monomers. We use a bond coefficient of $100.0\epsilon_{\mathrm{LJ}}$ and an equilibrium distance $r_{\mathrm{b}}=1.0\sigma$. For the Lennard-Jones interactions, we use a cut-off of $r_{\mathrm{c}}=2.5\sigma$ and attraction strength of $\epsilon_{\mathrm{LJ}}=1.0$. A strength of $\epsilon=1.0$ is also used for excluded volume interactions of the WCA potential. The charge of polymers is varied by simulation. The charge of ions is set to have a magnitude of $1.0$. Electrostatic interactions are calculated using a PPPM grid [50] for interactions beyond $2.5\sigma$ with relative error in forces of less than $10^{-4}$. Electrostatic interactions within $2.5\sigma$ are calculated directly. The dielectric constant is set to $\varepsilon=80.0$.
Each simulation trajectory is first annealed at a temperature of $k_{\mathrm{B}}T/\epsilon_{\mathrm{LJ}}=5.0$ for $10^{4}\tau$ using a Langevin thermostat with a damping of $100.0\tau$. Thereafter, the temperature of each simulation is set to $k_{\mathrm{B}}T/\epsilon_{\mathrm{LJ}}=1.0$ also using a Langevin thermostat with a damping of $100.0\tau$. Simulation trajectories are then run for $10^{6}\tau$. The last $5\times 10^{5}\tau$ is collected for data analysis, sampling a frame every $500\tau$. We use a timestep of $0.005\tau$ in the simulations. Additionally, all simulations in this study are performed in triplicate. Errors depicted in simulation figures correspond to the standard error between these simulation replicates during the sampling period of $5\times 10^{5}\tau$.
In the analysis of MD simulations, clusters (droplets) are defined using a cutoff of $2.0\sigma$ between polymers. The trajectory analysis software Ovito [51] is used to generate cluster data from simulation frames. An in-house code is used to generate radial density profiles from an Ovito [51] cluster analysis pipeline.
@@ A.2 Additional result figures
Additional figures are presented in this section corresponding to results from MD simulations discussed in the main text. Figure S1 presents radial charge density profiles from MD simulations at ($z_{P^{+}}=0.55$, $z_{P^{-}}=-0.55$), ($z_{P^{+}}=0.64$, $z_{P^{-}}=-0.3$), and ($z_{P^{+}}=1.05$, $z_{P^{-}}=-0.34$). See Main Figure 1b for a visualization of these droplets. Figure S2 shows normalized droplet size distributions corresponding to four charge pairings in Main Figure 4a. Figure S3 analyzes the relationship between charge asymmetry and droplet size across three different values of $\epsilon_{\mathrm{LJ}}$. Figure S4 depicts the standard error in average droplet sizes across simulation replicates. Finally, Figure S5 quantifies the standard deviation in normalized droplet sizes within the simulations presented in Main Figure 4a.
@@
Figure S1: Radial charge density profiles indicate charge asymmetry controls droplet size. Radial charge density profiles for (a) $z_{P^{+}}=0.55$, $z_{P^{-}}=-0.55$ (b) $z_{P^{+}}=0.64$, $z_{P^{-}}=-0.3$ and (c) $z_{P^{+}}=1.05$, $z_{P^{-}}=-0.34$. Shaded regions indicate the standard error from three independent simulation replicates. The profiles indicate that increasing charge asymmetry decreases droplet size. They also demonstrate that ions are excluded from the dense phase. See Main Figure 1b for visualization of these droplets.
@@
Figure S2: Charge asymmetries create distributions of droplet sizes. Four simulation systems, (I) $z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$ (II) $z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$ (III) $z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$ (IV) $z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$, corresponding to Main Figure 4a are plotted in histograms of normalized droplet sizes. Error bars indicate the standard error across histograms generated from three independent simulation replicates. At larger charge asymmetries (III and IV), a smaller average droplet size is favored. Finite-sized effects prevail at smaller charge asymmetries (I and II) indicated by a wide distribution of droplet sizes and high errors.
@@
Figure S3: Charge asymmetry’s effect on droplet size is preserved at higher polymer attraction strengths. The normalized droplet size for four charge pairings [($z_{P^{+}}=0.89$, $z_{P^{-}}=-0.54$), ($z_{P^{+}}=0.99$, $z_{P^{-}}=-0.5$), ($z_{P^{+}}=1.09$, $z_{P^{-}}=-0.46$), ($z_{P^{+}}=1.19$, $z_{P^{-}}=-0.42$) corresponding to I-IV in Main Figure 4a] at three different attraction strengths between oppositely charged polymers ($\epsilon_{\mathrm{LJ}}=1.0,1.25,1.5$). As charge asymmetry is increased, the normalized droplet size in the simulations decreases regardless of the value of $\epsilon_{\mathrm{LJ}}$. Error bars represent the standard error in average normalized droplet size for three independent simulation replicates.
Figure S4: Error on average simulation droplet sizes indicate finite-sized effects. The standard error in the average droplet size (across three simulation replicates) for the charge pairings plotted in Main Figure 4a is shown. At intermediate charge asymmetries (charge asymmetries closer to the diagonal that still present a patterned phase), the presence of finite-sized effects is demonstrated by sustained droplets of different sizes (yellow/gray points).